### v3

# Beyond-Labels: Advancing Open-Vocabulary Segmentation With Vision-Language Models

Abstract
Self-supervised learning can resolve numerous image or linguistic processing problems when effectively trained. This study investigated simple yet efficient methods for adapting previously learned foundation models for open-vocabulary semantic segmentation tasks. Our research proposed ”BeyondLabels,” a lightweight transformer-based fusion module that uses a handful of image segmentation data to fuse frozen image representations with language concepts. This strategy allows the model to successfully actualize enormous knowledge from pre-trained models without requiring extensive retraining, making the model data-efficient and scalable. Furthermore, we efficiently captured positional information in images using Fourier embeddings, thus improving the generalization across various image sizes, addressing one of the key limitations of previous methods. Extensive ablation tests were performed to investigate the important components of our proposed method; when tested against the common benchmark PASCAL-5i, it demonstrated superior performance despite being trained on frozen image and language characteristics.
Index Terms:
Beyond-Labels, open-vocabulary semantic segmentation, Fourier embeddings, PASCAL-5i.
I Introduction
Recent research on self-supervised learning has yielded promising results. Methods such as MoCo [6] and DINO [4] have demonstrated that powerful image representations can be learned without utilizing labeled data. Interestingly, even though these models were only trained on unlabeled images, they appear to naturally detect objects in images, allowing for fundamental semantic segmentation [17, 25, 29].
Another technique relies on learning a shared representation for images and text from big image-text pairs available on the internet. CLIP [20] and ALIGN [15], taught using simple contrastive learning techniques, have exhibited impressive capacities to perform new tasks without additional training, such as image classifying in a ”zero-shot” way. As computational power increases, more powerful models are anticipated to emerge, trained on larger datasets, and producing better image categorization results. It has opened a way towards the development of open-vocabulary segmentation, image retrieval, and multimodal reasoning within wide-ranging applications involving vision-language tasks. Larger and more powerful models are envisioned with increases in computational power and larger available datasets for further gains in both accuracy and robustness of image classification, along with other applications involving vision language. The above-mentioned continuous improvement illustrates the increased importance of contrastive learning in closing the gap in visual and textual understanding.
Semantic segmentation is a harder operation involving labeling each image pixel with a certain category. Recent models, such as FCN [24], U-Net [22], DeepLab [5], and DPT [21], have had tremendous success in training deep neural networks on huge datasets created for segmentation. Traditional techniques, however, rely on training classifiers for specific categories, which poses a significant constraint. These models can only predict labels for the categories on which they were trained, therefore, they cannot detect or segment objects from new, previously unseen categories during testing.
In our work, we investigate efficient techniques to connect the powerful pre-trained vision-language models to solve the challenge of open-vocabulary semantic segmentation successfully, ”the segmentation of objects from any category based on their textual names or descriptions.” To capture fine-grained spatial information, we first encode the input image using a frozen, self-supervised image model supplemented with Fourier embeddings, particularly across changing resolutions. Fourier embedding is used to extract high-level spatial/frequency features from image data to help the model learn better. Simultaneously, we encode category names using a frozen language model to ensure a strong representation of textual descriptions. The integrated embeddings from two alternative modalities allow for exact matching of picture regions and textual descriptions, resulting in seamless segmentation across many categories. The key ingredient of our approach is the reliance on Fourier embeddings [27], [13], [18] to encode positional information without resorting to the commonly used pre-trained position embeddings. This enhances the generalization across images and results in smooth and consistent spatial encoding. A transformer-based fusion module receives the combined features. Taking into account data from the other modality, this module uses self-attention to update the features. To create the segmentation mask, we calculate the cosine similarity between each pixel and the text characteristics in each category. The dot product of the L2 normalized image and language embeddings is used to accomplish this.
Among the contributions we provide, this includes a simple, lightweight, portable, transformer-based fusion module, Beyond-Labels, which stitches effectively a series of open-vocabulary pre-trained models. To encode the position, unlike typical usage, common pre-trained position embeddings will not be considered but instead are realized by Fourier embeddings, which bring great improvement in generalization among different images but with smooth, steady spatial encoding guaranteed. We also present ablation studies to verify our approach’s main components. This demonstrates strong zero-shot performance and competitive results on the PASCAL few-shot benchmark.
Figure 1: Our proposed framework consists of a text encoder, a multi-model embedding fusion module, a decoder, and the image encoder utilizing a pre-trained vision-language model with Fourier embeddings. The multimodal embedding fusion module combines the features the frozen encoders have extracted from the text and image data. Before segmenting the image features according to how closely they resemble text features, the decoder upscales them to their original size.
II Related Work
II-A Semantic Segmentation of Unseen Categories
Teaching a network to segment new, previously unknown categories remains difficult, as most existing approaches only function with a predefined set of categories, requiring the test and training categories to match. Zero-shot semantic segmentation attempts to address this issue by employing word embeddings (semantic representations of category names) to recognize new categories without further training data. There are two primary approaches:
  • Generative approaches, such as ZS3Net [3] and CaGNet [12], combine an image segmentation model with a mechanism for producing image features for new categories via word embeddings.
  • Discriminative algorithms, such as SPNet [30] and LSeg [8], map pixels and category names into a shared embedding space, which is then used to forecast the class of each pixel.
II-B Fourier Embedding for Spatial Encoding
Recently, Fourier embeddings have become appealing alternatives to traditional position encodings in those tasks that require strong spatial representation. Unlike learned positional embeddings, which will struggle hard to generalize outside the training distribution, Fourier-based embeddings provide a continuous and smooth encoding of spatial locations. Recently, it has been the case that Fourier embeddings turn out to be useful in improving generalization for implicit neural representations. For example, NeRF [19] proposed using Fourier feature mappings to help the positional encoding for 3D reconstruction tasks. Since then, this idea has been applied to image segmentation, where models benefit from the smoothness and periodicity of Fourier-transformed coordinates. In this context, self-supervised learning for Perceiver IO [14] uses Fourier embeddings to map high-dimensional inputs to a latent representation, with the best performance across various multimodal tasks. Correspondingly, in semantic segmentation, Fourier Network Representations [26] are utilized to encode better spatial dependencies, reducing large labeled datasets. Our approach thus replaces conventional learned embeddings with integrated Fourier embeddings of spatial information for smooth generalization across image domains. This goes in line with previous work showing that continuous and structured position encoding is helpful, both for vision and language models.
II-C Pre-trained Language and Vision Models
In computer vision and natural language processing, self-supervised learning has made substantial progress. For example, methods such as contrastive learning (MoCo [6]), metric learning (BYOL [11]), and self-distillation (DINO [4]) have emerged as effective tools for extracting characteristics for subsequent tasks. Language models, on the other hand, are commonly pre-trained with masked language modeling (BERT [7]) or next-token prediction (GPT [2]).
Large-scale image-language models have recently received increased interest. CLIP [20] is an important example, which trains on massive image-text pairs using contrastive learning and has successfully aligned images and text in a common space. Following that, CLIP’s pre-trained knowledge is applied to various tasks, including object detection, picture captioning, image segmentation, and text-driven image manipulation. Instead of fine-tuning CLIP for each job, we adopt a simpler strategy, integrating a lightweight fusion module with pre-trained self-supervised vision and language models. This methodology is an incredibly effective foundation for the semantic segmentation of new groups.
As foundation models [10, 9, 16] evolve rapidly, numerous studies have investigated how to merge multiple pre-trained models for different tasks. For example, Flamingo [1] proposed an architecture that uses big, pre-trained vision-only and language-only models to perform image-language tasks with minimal data. Socratic Models [31] uses a similar method but integrates pre-trained language models, vision-language models, and audio-language models to do multi-modal tasks like image captioning and video-to-text retrieval without retraining the models. Frozen [28] suggested a method for aligning a pre-trained, frozen language model with an image encoder using only a few examples and continual prompts.
III Methodology
This section begins by defining the open-vocabulary semantic segmentation problem. Next, we describe our suggested architecture, ”Beyond-Labels,” which links the previously trained language and image models to accomplish the task.
III-A Problem Formulation
In this section, we will describe our approach to formulating the data. A training set is provided to us:
$\text{Data}_{\text{train}}$ =$=\{(\mathcal{X},\mathcal{Y},\mathcal{U})\mid\mathcal{X}\in\mathbb{R}^{H\times W\times 3},$
$\mathcal{Y}\in\mathbb{R}^{H\times W\times|\mathcal{U}|},$
$\mathcal{U}\subseteq\mathcal{S}\}$ (1)
Where $\mathcal{X}$ is the input picture and $\mathcal{Y}$ is the one-hot encoded segmentation masks for $\mathcal{X}$; $\mathcal{U}$ represents a unit of observed categories in $\mathcal{X}$, and $\mathcal{S}$ represents all training categories in this unit. We aim to train a segmentation model that can separate a test image into significant areas of unexplored categories.
III-B Image Feature Extraction
The process starts with splitting the input image $\mathcal{X}$ as a collection of non-overlapping 2D patches. Each patch is flattened into a vector and put onto an embedding space with dimension ’d’ using a linear projection. Fourier embeddings [27] are used to encode positional information rather than standard pre-trained positional embeddings, allowing for improved generalization to images of varied sizes while also maintaining smooth and consistent spatial encoding. A Fourier embedding is calculated for each patch at point (x,y) on the grid using sinusoidal functions. The Fourier embedding is applied to the patch embedding, resulting in a final position-aware embedding $\mathcal{Z}$ for each patch.
$z_{i}$ $=x_{i}+f\_emb(x,y)$ (2)
Transformer encoder layers process position-aware embeddings, $\mathcal{Z}$ and layer normalization, feedforward networks, multi-head self-attention, and residual connections are used by each layer to analyze the embeddings. CLIP[20], a self-supervised language model trained on a huge collection of texts or images, is used as an image encoder.
$F_{V}$ $=\Phi_{\text{IE}}\left(f\_{\text{emb}}(X_{j})\right)\in\mathbb{R}^{(HW)\times d}$ (3)
Where ”d” is the image embedding dimension, p is the patch size, and H = H/p and W = W/p.
III-C Textual Feature Extraction
Semantic categories ($\mathcal{W}$) are encoded using a text encoder to produce text embeddings.
$F_{W}$ $=\Phi_{\text{TE}}(W)\in\mathbb{R}^{|W|\times d}$ (4)
Where ”d” is the dimension of the word embeddings and ”W” is the number of input categories. We employ prompt templates such as ”a picture of a category within the context” to provide semantic category context. The output embeddings generated by the text encoder are then averaged. The ”Result Section” offers a collection of prompt templates. We use CLIP [20], a self-supervised language model, as a text encoder.
Figure 2: Overview of the decoder section. Multimodal embedding fusion combines image and text encoder features that are processed through a decoder to generate upsampled image features. These were then compared with the textual features using cosine similarity, creating a segmentation mask with labels refined by thresholding and a temperature-scaled sigmoid.
III-D Multi_Model Embedding Fusion
To align the image and textual information, we use MLPs ($\theta$) to align their channel dimensions ($F_{V}$ and $F_{W}$). To capture the interactions between the textual and image data, we subsequently pass the combined features through a multimodal embedding fusion module:
$[F_{V}^{0},F_{W}^{0}]=\Phi_{\text{Fuse}}(\theta[F_{V},F_{W}])$ (5)
The fusion module uses multiple transformer encoder layers to capture image-text relationships. The fused image features ($F_{V}^{0}$) and text features ($F_{W}^{0}$) resemble the originals ($F_{V}$) and ($F_{W}$) in shape, but they are improved and enhanced by constantly focusing on one another’s modality. This enrichment comes because the fusion module iteratively puts more focus on cross-modal interactions, which allow the features of the image and text modalities to enrich each other’s representations. Thus, the fused features become better aligned and may be capable of providing a more holistic representation of the underlying data, therefore boosting performance in further tasks involving either images or text. Such an attention-driven fusion process significantly enhances the model’s capability of capturing fine-grained correlations between the two modalities, which is crucial for achieving high-quality results on tasks like open-vocabulary semantic segmentation. Our approach exploits the complementary strengths of both image and text data to allow more accurate and context-aware embeddings
III-E The Hierarchical Decoder
The multimodal embedding fusion combines the image and text encoders’ features. These combined features are passed to the decoder for further processing. The decoder gives us the image feature map, which is progressively upsampled to match the resolution of the input image. The upsampled image features are compared with the textual features using cosine similarity to generate the logits ŷ. This comparison helps determine how closely each area in the image aligns with the provided text categories. The final result is a segmentation mask, where different regions of the image are assigned labels. Using class-wise thresholding and a sigmoid with a temperature ($\mathcal{\tau}$), final predictions can be made using binary segmentation for each category.
IV Experimental Results
Model 5_0 5_1 5_2 5_3 mIoU
SPNet [30] 23.8 17.0 14.1 18.3 18.3
ZS3Net [3] 40.8 39.4 39.3 33.6 38.3
OSLSM [23] 33.6 55.3 40.9 33.5 40.8
Ours 44.6 49.4 40.0 32.1 41.5
TABLE I: Performances of different methods on the evaluation dataset of PASCAL-5i.
Model F_Emb Fusion Image_Decoder 50 51 52 53 mIoU
B_L_0 X X 32.6 34.1 25.2 21.0 25.7
B_L_1 X 38.4 44.6 35.4 29.3 36.9
B_L_2 44.6 49.4 40.0 32.1 41.5
TABLE II: Ablation study on PASCAL-5i
IV-A Dataset
We evaluated our approach on the PASCAL-5i [23] dataset, which is designed for few-shot segmentation approaches. It is based on the well-known PASCAL VOC 2012 dataset. The dataset is divided into four folds (i = 0, 1, 2, 3) and contains 20 different object classes. Each fold contained five classes, and the model was trained on the other fifteen classes. This cross-validation design separates the classes used for testing from the classes used for training, ensuring that this is indeed a valid evaluation strategy to determine how well few-shot learning algorithms generalize onto new categories. The support set contains some labeled instances, for example, 1-shot or 5-shot, from each of the five target classes in each fold to use as a reference for segmentation. Given this limited information in the support set, the query set should separate the objects precisely in the unlabeled photos that contain objects from the same five classes. Models trained on 15 classes and tested on the remaining 5 are doing so to cover all classes iteratively, using a technique called ”leave-one-fold-out.” Since this is indeed an exhaustive evaluation framework, it has made PASCAL-5i a standard benchmark for the comparison of few-shot segmentation. We have reported the results of our method in the ablation study and results comparison section.
IV-B Evaluation Metric
Our key evaluation metric is mean intersection-over-union (mIoU), which averages IoU across all classes within the fold.
$\text{mIoU}=\frac{1}{F}\sum_{F=1}^{F}\text{IoU}_{F}$
Where IoU denotes the intersection over the union of class (F), which is the target fold’s class count.
IV-C Results Comparison
Table 2 compares different models on the segmentation task for all folds in terms of the overall mean IoU, namely, mIoU. SPNet has a representation with an overall mean IoU of 18.3 and folds 2 and 3 performed the weakest, showing the low ability of the network to generalize well against the unseen complexities of classes within few-shot segmentation tasks. ZS3Net performed better with a mIoU of 38.3, trying to generalize across folds in semantic features The OSLSM improves further by achieving a mIoU of 40.8, which is superior to that of ZS3Net and SPNet; however, challenges remain for particular folds such as 3. Our model surpasses all baselines by a mIoU of 41.5, with maximum performance at folds 0 and 1 at 44.6 and 49.4, respectively. This again supports the hypothesis that integrating Fourier embedding with multi-model fusion, together with a decoder, will improve feature representation by paying full consideration to the mutual image and textual information. This indicates that the relative improvement of 3.2 in mIoU over OSLSM and 8.4 over ZS3Net demonstrates the effectiveness of our proposed components in handling complex spatial relationships and unseen class distributions. These results confirm that our approach not only achieves better segmentation accuracy but also ensures more consistent performance across diverse folds.
IV-D Ablation Study
We investigated the contribution of each core component in the proposed ”Beyond-Labels” framework. Specifically, we investigated the impact of the Fourier Embedding (F_Emb), multi-model embedding fusion module (Fusion), and the use of a decoder instead of traditional positional embedding techniques. Detailed results of this study can be seen in Table 1 for breaking down performance in different configurations.
  • By default, it uses only an upward model that includes the usual positional embedding and upsampling using a decoder. The use of this standard baseline model resulted in minimum performance, by which B_L_0 obtained a mIoU of only 25.7. The large decreases in scores in folds 2 and 3 represent a limitation based on the use of standard position embedding without robust feature fusion in action. Without Fourier embedding, the model cannot capture such high-dimensional and expressive positional information, which is critical for improving segmentation accuracy, especially in challenging folds such as folds 2 and 3.
  • While for B_L_1, incorporation of Fourier Embedding improves its performance on all counts with the mIoU increasing to 36.9. This, in turn, indicates that adding frequency-based representation is highly vital for boosting the feature extraction required for improving the generalization of non-visible classes. More important is the model’s capability of adaptation to different complicated spatial scenarios due to improved performances on all folds.
  • The best-performing model, B_L_2, is the full model, including the Fourier embedding, Fusion, and decoder, with a mIoU of 41.5. Additional positional information is introduced by Fourier Embedding that, combined with the fusion module, enables the model to learn richer representations. Further refinement of the upsampling process with the decoder leads to consistent improvements in all the folds. This includes the Fourier embedding module with the Fusion module, which generates high-dimensional and expressive forms of position-based information, thus yielding significant gains in segmentation accuracy. This further facilitates better spatial understanding within the image, boosting the capability for more precise feature representations of the model.
The benefits of Fourier embedding are easily found in all folds, where traditional embeddings often fall short. This can be integrated into the full model to demonstrate how higher-quality positional encoding can be combined with other components to achieve top-level performance.
IV-E Implementation Details
The vision-language model (CLIP [20]) is frozen throughout the training process. Each of the 8 levels that make up the multi-model embedding fusion module has 4 attention heads. After k convolutional layers, the decoder performs a 2× upsampling operation. We employ the AdamW optimizer with a cosine scheduler to decay the learning rate to maximize training. We use cross-entropy for the loss function and set the temperature factor at ($\mathcal{\tau}$)= 0.08.
IV-F Conclusion
Our research aims to make advanced vision and language models more helpful for open-vocabulary semantic segmentation. We introduce ”Beyond-Labels,” a compact and efficient module that links image and language models to do open-vocabulary semantic segmentation, which involves finding and categorizing items in images, even for previously unknown categories. We tested it on the “PASCAL” dataset and found that it performs extremely well, exceeding some existing approaches. Despite its simplicity, “Beyond-Labels” delivers excellent results.
IV-G Future Work
To assess our work’s resilience and enhance generalization across many domains, we will expand it to include more datasets and perform more ablation studies. Through testing on various datasets, we hope to ensure that our method works reliably and successfully in practical situations, capturing a wider range of difficulties and subtleties in semantic segmentation tasks.
IV-H Template List
  • An image of a {category}.
  • This is an image of a {category}.
  • An image of a small {category}.
  • An image of a medium {category}.
  • An image of a large {category}.
  • An image of a {category} within the context.
  • An image of the {category} within the context.
  • An image of the {category} within the context.
  • A resized image of a{category} within the context.
  • This falls under a {category} within the context.
  • This falls under the {category} within the context.
  • This falls under one {category} within the context.